A burglar coughs with a sound intensity of 2.48 * 10 ^ - 7 * W / (m ^ 2) . A burglar alarm will ring if sound is detected 30dB above the alarm's threshold of 1.5 * 10 ^ - 10 * W / (m ^ 2) Will the burglar set off the alarm ? 


1
Expert's answer
2021-01-26T13:39:19-0500

The sound intensity in decibel can be found as follows:


"I(dB)=10log_{10}\dfrac{I}{I_0},""I(dB)=10log_{10}\dfrac{2.48\cdot10^{-7}\ \dfrac{W}{m^2}}{1.5\cdot10^{-10}\ \dfrac{W}{m^2}}=32\ dB."

As we can see from the calculations, the burglar will set off the alarm.